OpenStudy (anonymous):

Factor out the greatest common factor. 16y^5 - 24y^4 - 40y^3

OpenStudy (anonymous):

Well, you can first factor out a y^3. Then you need to find the GCF of 16,24,40 16=2*8 24=3*8 40=5*8 So the GCF is 8y^3